OpenStudy (anonymous):

Which one is a rule for the nth term of the sequence 3, 15, 75, 375...?

OpenStudy (cwrw238):

note that each term is obtained from previous by multiplying by 5

OpenStudy (anonymous):

A. an = 3(5)n − 1 B. an = 5(3)n − 1 C. an = 3(5)n + 1 D. an = 5(3)n + 1

OpenStudy (cwrw238):

this is called a geometric sequence

OpenStudy (anonymous):

those are my answer choices

OpenStudy (cwrw238):

try plugging n = 2, 3 and 4 into A and see if you get 15,75 and 375

OpenStudy (cwrw238):

eg n=2: 3(5)^(2-1) = 3*5^1 = 3*5 = 15 - that one works now try n=3

OpenStudy (cwrw238):

do you see what i'm doing?